Q6. The set B = {1+t2, t+t, 1+2t+t2} is a basis for P2. Find the coordinate vector of p(t) = 3+t-6t2 relative to B.
Answer #1

Let wordinates a bit) =3+1-66² relative to basis B = s={1+th, tits, Itzt+t2} be Therefore, [] =utt-664 act+²) tblt+62)+((1+2t